OpenStudy (anonymous):

Find the derivative of f(x)=e^2x-5(X+3)

OpenStudy (anonymous):

find the derivative of \[f(x)=e^{2x-5}(x+3)\]

zepdrix (zepdrix):

So we have the product rule here :D Are you familiar with that?

OpenStudy (anonymous):

Hey Zepdrix, I think so yes

OpenStudy (anonymous):

what I have so far is (e^2x-5)(x+3) + (e^2x-5)(1)= xe^2x-5 + 3e^2x-5+e^2x-5 = xe^2x-5 + 4e^2x-5= e^2x-5(x+4)

zepdrix (zepdrix):

Your derivative of the exponential term in the first part has a small mistake. When you take the derivative of e^(something) Yes, you get back e^(something).. but you also need to apply the chain rule, You will end up with (derivative of something)*e^(something)

OpenStudy (anonymous):

OK...Thanks

zepdrix (zepdrix):

Uh oh you sound confused XD

zepdrix (zepdrix):

 Does this help maybe? :O you were very very close